# Stable soap bubble clusters with multiple torus bubbles

### v6

@@ 3.1 Surface Evolver configuration
Figure 2: $2$D view of the cluster.
@@ 4.1.1 Surface Evolver configuration
A smoothed $2$D schematic view of the configuration might help to get a better understanding.
Figure 7: Smoothed $2$D schematic view of the cluster.
@@
The genus of a compact, closed, connected, orientable topological surface without boundary is defined as the largest number of simple closed curves with no common points that can cut the surface without making it disconnected. Simply put, the genus is the “number of holes” in the surface (in the sense of torus-like holes). However, consider now a solid whose boundary has genus 00 (sphere). Dig a cavity in the solid, with no contact with the exterior, and $n$ holes from outside to the cavity. Then the surface of the new solid has genus $n-1$, i.e. is topologically an $(n-1)$-fold torus. Roughly speaking, one can imagine that one of the holes is stretched until the surface can be placed on a plane, resulting in $n-1$ holes only. This can be illustrated in the following way for an excavated tetrahedron whose surface is homeomorphic to the bubble of figure 6.
@@
Figure 10: $2$D scheme of the configuration.
@@ 4.2.1 Surface Evolver configuration
The cluster is parameterized as in section 4.1.1 with the outer cube $[-1/2,1/2]^{3}$. To avoid possible instabilities, the scaling factors have been chosen to avoid large inner double bubbles. They have the following values.
@@ 4.3.2 Energy minimizing cluster
Figure 23: Left discretization: $1.6\times 10^{-3}$. Right discretization: $78\times 10^{-3}$.

### v7

@@ 3.1 Surface Evolver configuration
Figure 2: 22D view of the cluster.
@@ 4.1.1 Surface Evolver configuration
A smoothed 22D schematic view of the configuration might help to get a better understanding.
Figure 7: Smoothed 22D schematic view of the cluster.
@@
The genus of a compact, closed, connected, orientable topological surface without boundary is defined as the largest number of simple closed curves with no common points that can cut the surface without making it disconnected. Simply put, the genus is the “number of holes” in the surface (in the sense of torus-like holes). However, consider now a solid whose boundary has genus 0 (sphere). Dig a cavity in the solid, with no contact with the exterior, and $n$ holes from outside to the cavity. Then the surface of the new solid has genus n1n-1, i.e. is topologically an (n1)(n-1)-fold torus. Roughly speaking, one can imagine that one of the holes is stretched until the surface can be placed on a plane, resulting in n1n-1 holes only. This can be illustrated in the following way for an excavated tetrahedron whose surface is homeomorphic to the bubble of figure 6.
@@
Figure 10: 22D scheme of the configuration.
@@ 4.2.1 Surface Evolver configuration
The cluster is parameterized as in section 4.1.1 with the outer cube $1/2,1/2]^{3}$. To avoid possible instabilities, the scaling factors have been chosen to avoid large inner double bubbles. They have the following values.